Assume that when adults with smartphones are randomly​ selected, ​64% use them in meetings or classes. If 9 adult smartphone users are randomly​ selected, find the probability that exactly 2 of them use their smartphones in meetings or classes.


1
Expert's answer
2021-10-13T17:37:05-0400

Solution:

"p = 64 \% = 0.64"

n = 9

"X\sim Bin(n,p)"

"P(X=x) = \frac{n!}{x!(n-x)!} \times p^x \times (1-p)^{n-x} \\

P(X=2) = \frac{9!}{2!(9-2)!} \times 0.64^2 \times (1-0.64)^{9-2} \\
= 0.01155"